OpenStudy (anonymous):

Multiply. (5x)(3x)(6x^2 – 1) Answer 48x^3 – 8x 48x^4 – 8x^2 90x^3 – 15x 90x^4 – 15x^2

OpenStudy (anonymous):

90x^4 – 15x^2

OpenStudy (campbell_st):

start with 5x and 3x it gives 15x^2 now multiply each term inside the brackets by 15x^2

OpenStudy (anonymous):

do the same as instructed before multiply together the outer terms and then distribute each into parenthess.

OpenStudy (anonymous):

Yes as stated above: Use the following order for such equations: BODMAS B Rackets first O OF ( if there is of) D Division ( if there is division) M Multiplication ( if multiplication exists) A Additition ( if there is addition) S Subtraction ( if there is subtraction) .example: (5x)(3x)(6x^2 – 1) 1. work the brackets first 5x * 3x = 15x^2 2. Multiply the above into the left brackets: 15x^2 ( 6x^2 -1) = 15x^2*6x^2 - 15x^2*1 = 90x^4 - 15x^2\[90x ^{4} - 15x ^{2}\]